The period T (time in seconds for one complete cycle) of a simple pendulum is related to the length L (in feet) of the pendulum by the formula 8t^2=pi^2 L. If a child is on a swing with a 10-foot chain, then how long does it take to complete one cycle of the swing?
Answer by checkley77(12569)

8t^2=pi^2L
8t^2=3.14^2*10
8t^2=9.8596*10
8t^2=98.596
t^2=98.596/8
t^2=12.3245
t=sqrt12.3245
t=3.51 sec.